News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Friday 9 August 2013 9.00am EDT

During the last week Chandra completed the observing schedule as planned.

Chandra passed through the sixth, seventh, and eighth eclipses of the season on Aug 3, 6, and 8 respectively, with nominal power and thermal performance.

A real-time procedure was executed on Aug 3 to perform a check-out of the B-side MUPS thrusters, which had not been fired since Aug 1999. The check-out goals were to assess the safe-operation and performance of the B-side thrusters prior to moving forward with a swap to using them for routine momentum unloads, saving the A-side thrusters for safe-mode operation. The two thrusters on the cold side of the spacecraft exhibited nominal thrust levels; however, the two on the hot side showed reduced thrust. The under-performance was anticipated as a possibility by the engineering team and it is theorized that it is due to gaseous helium bubbles passing through the thrusters. It is expected that these bubbles will pass with additional thruster pulses.

A real-time procedure was executed on Aug 7 to perform a routine self-check of the Electrical Interface Assembly (EIA) Sequencer.

Of note this week was the release on Aug 6 of CXCDS 10.0.1 for standard processing, which includes a small number of bug-fixes and enhancements, and on Aug 7 of version 4.5.4 of the CIAO Script package.

The schedule of targets for the next week is shown below and includes an observation of the Crab coordinated with HST and an observation of Sgr A coordinated with the EVLA.

        SgrA                   ACIS-S       Aug 11
        SDSSJ161027.41+1308    ACIS-S       Aug 12
        IRDCG34.4+0.23IRD      ACIS-I
        G113.82+44.35          ACIS-I       Aug 13
        3C455                  ACIS-S
        G139.59+24.18          ACIS-I
        A665                   ACIS-I
        Radiation Belts
        HE0435-1223            ACIS-S       Aug 14
        Abell85                ACIS-I
        CAN010                 ACIS-I       Aug 15
        WR6                    ACIS-S/HETG
        Radiation Belts                     Aug 16
        Crab                   ACIS-S
        SN2012hn               ACIS-S       Aug 17
        Abell85                ACIS-I
        Abell2443              ACIS-I
        A7                     ACIS-I       Aug 19


All spacecraft subsystems continued to support nominal operations.

Return to Status Reports